A test for drug use involves addition of 5.00mL of colorless reagent to 10.00 mL urine.If drug residue is present in the sample the solution turns blue(λmax = 605 nm, ε = 18,200M-1cm-1). Because other things in urine may give slightly positive results, only tests that measure a minimum concentration of 5.0 x 10-5 M drug residue in the original urine sample are considered positive for illegal drugs. When you measure the sample using the test described above and get an Abs of 0.28 at 605 nm (b = 1 cm). Can you conclude that this person has been using illegal drugs? Explain.

Show your work please.
Answer #1

The relation between absorption (A), Concentration (C), Path length of sample (b) and Molar extinction coefficient () can be given as below

A = 0.28 for sample, = 18200 M-1 cm-1 , b = 1 cm

C= 0.28 / 18200 x1 =1.5348 x 10 -5

To conclude that a person is using drug, the minimum concentration of drug should be 5 x 10 -5 M. But, from the absorption data the concentration found is less (1.5348 x 10 -5 < 5 x 10 -5 M).

So, we can conclude that the person is not using illegal drugs in the recent times.